# Supersingular curves via the Shimura–Taniyama method

### v1

@@ Abstract.
For a curve which admits an abelian cover of the projective line branched at three points, we study when its reduction to positive characteristic is supersingular. Using the method of Shimura and Taniyama, we give a complete classification when the genus of the curve is at most 10. The natural density of the set of primes for which this construction yields a supersingular curve is larger than expected.
@@
Pries was supported by NSF grant DMS-22-00418. We thank the organizers of the RIMS conference on Theory and Applications of Supersingular Curves and Supersingular Abelian Varieties, II. We thank Everett Howe, Wanlin Li, and Elena Mantovan for helpful conversations.
@@ 1. Introduction
Suppose $k$ is an algebraically closed field of prime characteristic $p$. An elliptic curve over $k$ is supersingular if it has no points of order $p$. A smooth projective curve $X$ of genus $g$ over $k$ is supersingular if its Jacobian is isogenous to a product of $g$ supersingular elliptic curves.
One good method for constructing supersingular curves is to study a curve $X$ that is an abelian cover of the projective line $\mathbf{P}^{1}$ branched at exactly three points. The degree and ramification indices of the cover determine the genus $g$ of $X$. Such curves are quotients of a Fermat curve and their Jacobians have complex multiplication. This makes it easier to study the action of Frobenius on the cohomology [Yui80] and hence determine whether the reduction of $X$ modulo $p$ is supersingular.
Recently, we were curious which pairs $(g,p)$ occur for supersingular curves in this context. Our motivation was to show that a new construction of supersingular curves of genus $5$ [BP] applied for infinitely many new primes. We wrote this paper because we could not find a reference that states what can be determined using this method.
Let $\mathbf{S}_{g}$ be the set of primes $p$ for which there exists a supersingular curve $X$ of genus $g$ defined over $k$ such that $X$ is an abelian cover of $\mathbf{P}^{1}$ branched at exactly three points. In this paper, we determine $\mathbf{S}_{g}$ for $5\leq g\leq 10$. The main result is Theorem 6.1, which builds on data found in Sections 3, 4 and 5. We summarize it as follows.
@@
For $5\leq g\leq 10$, here is the natural density $\delta_{g}$ of the set $\mathbf{S}_{g}$:
@@
The cases $g=9,10$ are especially interesting. The codimension of the supersingular locus in the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties of dimension $g$ is $g(g+1)/2-\lfloor g^{2}/4\rfloor$. For $g\geq 9$, this exceeds the dimension of the moduli space $\mathcal{M}_{g}$ of smooth curves of genus $g$. For this reason, the existence of a smooth supersingular curve of genus $g$ is geometrically unexpected for $g\geq 9$.
Work of Yui [Yui80] implies that $\delta_{g}$ is positive for all $g$, which can also be deduced from our analysis, see Remark 5.3.
For the proof, we use Ekedahl’s work on [Eke87, page 173] to classify the degrees and ramification indices of abelian covers $X\to\mathbf{P}^{1}$ branched at 3 points with fixed genus $g$. When $g=4,5$, Ekedahl determines conditions on $p$ for which $X$ is superspecial (meaning its Jacobian is isomorphic to a product of supersingular elliptic curves). We work instead with the Shimura–Taniyama formula [Tat71, Section 5], a method used also in [LMPT19b], to determine the set of primes $p$ for which the reduction is supersingular. We use SageMath to enumerate all the possibilities for genus $5\leq g\leq 10$.
We restrict to the case $g\geq 5$ and $p$ odd because of earlier work: for $g=1$ by Deuring [Deu41]; for $g=2$ by Serre [Ser83, Théorème 3], also [IKO86, Proposition 3.1]; for $g=3$ by Oort [Oor91, Theorem 5.12], also [Ibu93, Theorem 1]; and for $g=4$ by Harashita, Kudo, and Senda [KHS20, Corollary 1.2], also [Pri24, Theorem 1.1]. For $p=2$, for all $g$, see the work of van der Geer and van der Vlugt [vdGvdV95].
@@ 2. Abelian covers branched at three points
2.1. The degree and ramification indices
@@
Any two of $e_{0,1}$, $e_{0,\infty}$, and $e_{1,\infty}$ are relatively prime and $e_{0,1}\geq e_{0,\infty}\geq e_{1,\infty}$.
@@
(2.1) $2g-2=s(re_{0,1}e_{0,\infty}-(1+(e_{0,1}/e_{1,\infty})+(e_{0,\infty}/e_{1,\infty})).$
@@
The result follows since $c_{0}=re_{0,1}e_{0,\infty}$, $c_{1}=re_{0,1}e_{1,\infty}$, and $c_{\infty}=re_{0,\infty}e_{1,\infty}$. ∎
We identify $G$ with $\mathrm{Gal}(\phi)$ and define a refinement of the ramification type.
@@
See e.g., [Bou01, Definitions 3.2 & 4.1]. The inertia type $a$ is the triple $a=(a_{0},a_{1},a_{\infty})$, where $a_{b}$ denotes the canonical generator for $I_{b}$, found from the action of the inertia group on a local parameter at a ramification point above $b$.
There are three constraints on the inertia type: $|a_{b}|=c_{b}$ for $b\in B$; $a_{0},a_{1},a_{\infty}$ generate $G$; and $a_{0}a_{1}a_{\infty}=\mathrm{Id}_{G}$. This implies that $s$ divides $c_{1}$ and $s$ divides $c_{\infty}$. Not every ramification type $z=[c_{0},c_{1},c_{\infty},s]$ has an inertia type that satisfies these constraints. See Remark 5.1 for some examples.
Given $z$, we can change the labeling of branch points whose inertia groups have the same order; this allows us to permute these entries of the inertia type. Also, we can change the identification of $G$ with $\mathrm{Gal}(\phi)$; this modifies each entry of the inertia type by the action of some $\sigma\in\mathrm{Aut}(G)$. Two inertia types are equivalent, denoted $\approx$, if they are in the same orbit under these actions.
@@ 2.2. Cyclic covers branched at three points
In this section, we suppose $\phi$ is cyclic of order $m$, and identify $G=\mathbf{Z}/m\mathbf{Z}$. The inertia type $(a_{0},a_{1},a_{\infty})$ is defined up to multiplication by $(\mathbf{Z}/m\mathbf{Z})^{*}$. Note that $c_{b}=|a_{b}|=m/\gcd(m,a_{b})$. An affine equation for the cover is
$X:y^{m}=x^{a_{0}}(1-x)^{a_{1}},$
@@ 2.3. Non-cyclic Galois groups and inertia types
The following determines the Galois group in all the cases with $s>1$ for $5\leq g\leq 10$.
@@
By hypothesis, there is a prime $\ell$ dividing all three of $s=c_{0}s/c_{0}$, $c_{0}s/c_{1}$, and $c_{0}s/c_{\infty}$. If $G$ was cyclic, $\ell$ would divide each of $m,a_{0},a_{1},a_{\infty}$, contradicting the fact that $a_{0},a_{1},a_{\infty}$ generate $G$. Since $G$ is generated by two elements $a_{0},a_{1}$, this implies that $G\simeq\mathbf{Z}/r_{1}\mathbf{Z}\times\mathbf{Z}/r_{2}\mathbf{Z}$ where $r_{1}r_{2}=c_{0}s$ and $r_{2}\mid r_{1}$. Note that $r_{1}$ is a multiple of $c_{0}$ since $a_{0}$ has order $c_{0}$. The hypothesis on $z$ implies that the orders of $a_{1}$ and $a_{\infty}$ divide $c_{0}$. So $a_{0}$ and $a_{1}$ generate $G$ only if $r_{1}=c_{0}$. Thus $G\simeq\mathbf{Z}/c_{0}\mathbf{Z}\times\mathbf{Z}/s\mathbf{Z}$. ∎
For $s>1$, we denote elements of $G=\mathbf{Z}/c_{0}\mathbf{Z}\times\mathbf{Z}/s\mathbf{Z}$ as an ordered pair. So each $a_{b}$ in the inertia type $a=(a_{0},a_{1},a_{\infty})$ is written as $a_{b}=(a_{b,1},a_{b,2})$. To choose a representative for the equivalence class of $a$, we take the natural inclusion of $G$ in $V=(\mathbf{Z}/c_{0}\mathbf{Z})^{2}$, and use that $\mathrm{Aut}(V)$ acts transitively on pairs of linearly independent elements of order $c_{0}$. Without loss of generality, we choose $a_{0}=(1,0)$. When $c_{\infty}=s$, we also choose $a_{\infty}=(0,1)$.
2.4. Newton polygons for cyclic covers branched at three points
Suppose $p$ is a prime not dividing $m$, the degree of the cyclic cover $\phi$. We consider the reductions of $X$ and ${\mathbf{Jac}}(X)$ modulo $p$. The Newton polygon of ${\mathbf{Jac}}(X)$ is an invariant that measures the $p$-adic valuations of the eigenvalues of Frobenius. In this section, we explain how to compute the Newton polygon from the signature type of $\phi$.
Definition 2.7.
Suppose $A$ is a principally polarized abelian variety of dimension $g$ over $k$, By the Dieudonné–Manin classification, the $p$-divisible group $A[p^{\infty}]$ is isogenous to $\oplus_{\lambda}G_{\lambda}^{m_{\lambda}}$ for some multi-set of non-negative integers $m_{\lambda}$. This multi-set defines a symmetric Newton polygon $\nu(A)$ of height $2g$ with integer break-points.
Remark 2.8.
Suppose $A$ is defined over $\mathbf{F}_{0}$, a finite field of characteristic $p$. Let $W(\mathbf{F}_{0})$ denote the Witt vector ring of $\mathbf{F}_{0}$. Consider the action of Frobenius $\varphi$ on the crystalline cohomology group $H^{1}_{\rm cris}(A/W(\mathbf{F}_{0}))$. There exists an integer $n$ such that $\varphi^{n}$, the composite of $n$ Frobenius actions, is a linear map. The Newton polygon $\nu(A)$ is the multi-set of rational numbers $\lambda$ such that $n\lambda$ are the valuations at $p$ of the eigenvalues of Frobenius for this action.
We sometimes describe the Newton polygon of a curve $X$ in terms of the decomposition of ${\mathbf{Jac}}(X)[p^{\infty}]$. For example, a curve of genus $g$ is ordinary if its $p$-divisible group is $(G_{0,1}\oplus G_{1,0})^{g}$ and is supersingular if its $p$-divisible group is $G_{1,1}^{g}$.
Notation 2.9.
Let $R$ denote the set of orbits of the multiplication by $p$ map [×$[\times p]$ on the set $\{j\in\mathbf{Z}/m\mathbf{Z}:j\neq 0\}$. For each orbit $r\in R$, the additive order $d_{j}$ of $j$ modulo $m$ is the same for all values $j$ in $r$ and we denote it by $d_{r}$. For $\epsilon\in\{0,1\}$, let
[…]
The set $R$ is in one-to-one correspondence with the set of primes of the group ring $\mathbf{Q}[\mu_{m}]$ above $p$. The $p$-divisible group ${\mathbf{Jac}}(X)[p^{\infty}]$ canonically decomposes into pieces ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ indexed by the orbits $r\in R$.
By the Shimura–Taniyama formula [Tat71, Section 5], the only slope of the Newton polygon of ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ is $\lambda_{r}:=\alpha_{r}/\#r$. In particular, if every orbit is self-dual, meaning invariant under negation modulo $m$, then $X$ is supersingular.
Example 2.10.
If $a=(1,1,6)$, then $\phi$ has signature $f=(1,1,1,0,0,0,0)$. Thus $S_{1}=\{1,2,3\}$. Then $\lambda_{r_{1}}=1$, $\lambda_{r_{2}}=1/2$, and $\lambda{r_{5}}=0$. Thus $\nu(\mathbf{Jac}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$.
Remark 2.11.
One can also determine whether $X$ is supersingular using a method of Yui [Yui80] and Aoki [Aok08] that relies on $p$-adic valuations of Jacobi sums. By [Aok08, Theorem 1.1], $X$ is supersingular if and only if either (i) $p^{i}\equiv-1\bmod m$ for some $i$; or (ii) $a\approx(1,-p^{i},p^{i}-1)$ for some $i$ such that $d=\mathrm{gcd}(p^{i}-1,m)>1$ and $p^{j}\equiv-1\bmod m/d$ for some $j$.

### v2

@@ Abstract.
For a curve $X$ which is an abelian cover of the projective line branched at three points, we study when its reduction to positive characteristic is supersingular. Using the method of Shimura and Taniyama, we give a complete classification of primes $p$ for which the reduction of $X$ modulo $p$ is supersingular when the genus of $X$ is at most $10$. The natural density of this set of primes is larger than expected.
@@
Key words and phrases:
curve, Jacobian, abelian variety, positive characteristic, supersingular, Frobenius, Newton polygon, automorphism
Pries was supported by NSF grant DMS-22-00418. This work was supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University We thank the organizers of the RIMS conference on Theory and Applications of Supersingular Curves and Supersingular Abelian Varieties, II. We thank Everett Howe, Wanlin Li, Elena Mantovan, and Darren Schmidt for helpful conversations, and the referee for helpful comments.
@@ 1. Introduction
Let $k$ be an algebraically closed field of prime characteristic $p$. An elliptic curve over $k$ is supersingular if it has no points of order $p$. A smooth projective curve $X$ of genus $g$ over $k$ is supersingular if its Jacobian ${\mathbf{Jac}}(X)$ is isogenous to a product of $g$ supersingular elliptic curves.
One good method for constructing supersingular curves is to study families of curves that are abelian covers of the projective line $\mathbf{P}^{1}$. The degree and ramification indices of the cover determine the genus $g$. When the curve $X$ is branched at exactly three points, it is a quotient of a Fermat curve and ${\mathbf{Jac}}(X)$ has complex multiplication. This makes it easier to study the action of Frobenius on the cohomology [Yui80] and hence determine whether the reduction of $X$ modulo $p$ is supersingular. For example, Re studied supersingular quotients of the Fermat curve with the aim of producing supersingular curves of many genera in fixed characteristic [Re08, Re09].
Recently, we were curious about an analogous question: given $g$, for which primes $p$ do these constructions yield a supersingular curve of genus $g$ in characteristic $p$? Our motivation was to show that a new construction of supersingular curves [BP] yields supersingular curves of genus $5$ in characteristic $p$ for infinitely many new primes $p$. We wrote this paper because we could not find a reference that states what can be determined using this method.
Let $\mathbf{S}_{g}$ be the set of primes $p$ for which there exists a supersingular curve $X$ of genus $g$ defined over $k=\overline{\mathbf{F}}_{p}$ such that $X$ is an abelian cover of $\mathbf{P}^{1}$ branched at exactly three points. For $5\leq g\leq 10$, we determine the primes for which there exists a supersingular curve of this type in Sections 3, 4 and 5. This yields the main result of this paper, Theorem 6.1, which states exactly which primes are in $\mathbf{S}_{g}$, for $5\leq g\leq 10$, in terms of congruence conditions. From these congruence conditions, we determine the natural density of the set $\mathbf{S}_{g}$, which we state as the following corollary.
@@
For $5\leq g\leq 10$, the natural density $\delta_{g}$ of the set $\mathbf{S}_{g}$ is:
@@
[…]
Work of Yui [Yui80] implies that $\delta_{g}$ is positive for all $g\geq 1$, which can also be deduced from our analysis, see Remark 5.3.
For the proof, we use Ekedahl’s work [Eke87, p. 173] to classify the degrees and ramification indices of abelian covers $X\to\mathbf{P}^{1}$ branched at $3$ points, such that $X$ has genus $g$. When $g=4,5$, Ekedahl determines conditions on $p$ for which $X$ is superspecial (meaning ${\mathbf{Jac}}(X)$ is isomorphic to a product of supersingular elliptic curves). We determine the set of primes $p$ for which the reduction is supersingular using the Shimura–Taniyama formula [Tat71, Section 5], a method also used in [LMPT19b]. As explained in Section 2.4, this method exploits the structure of complex multiplication on ${\mathbf{Jac}}(X)$ to determine its Newton polygon.
We use SageMath to enumerate all abelian covers of $\mathbf{P}^{1}$ ramified at three points with genus $g$ satisfying $5\leq g\leq 10$. We restrict to the case $g\geq 5$ and $p$ odd because of previous constructions of supersingular curves: for $g=1$ by Deuring [Deu41]; for $g=2$ by Serre [Ser83, Théorème 3], also [IKO86, Proposition 3.1]; for $g=3$ by Oort [Oor91, Theorem 5.12], also [Ibu93, Theorem 1]; for $g=4$ by Harashita, Kudo, and Senda [KHS20, Corollary 1.2], also [Pri, Theorem 1.1]; and for $p=2$ and arbitrary $g$, by van der Geer and van der Vlugt [vdGvdV95].
Remark 1.2.
Many mathematicians studied supersingular curves of genus $g\geq 5$ in characteristic $p$ for various pairs $(g,p)$. Here is an incomplete list: Hermitian curves [Han92], Artin-Schreier curves [vdGvdV92, vdGvdV95, BHM+16], cyclic covers of $\mathbf{P}^{1}$ [Re08, Re09, LMPT19b, LMPT19a]; and trigonal curves [Kud19, KH22].
@@ 2. Abelian covers branched at three points
2.1. Ramification and Inertia Types
@@
Ekedahl [Eke87, page 173] verified that $e_{0,1}$, $e_{0,\infty}$, and $e_{1,\infty}$ are pairwise relatively prime, that $e_{0,1}\geq e_{0,\infty}\geq e_{1,\infty}$; and that
(1) $c_{0}=re_{0,1}e_{0,\infty},\quad c_{1}=re_{0,1}e_{1,\infty},\quad c_{\infty}=re_{0,\infty}e_{1,\infty}.$
@@
(2) $2g-2=s(re_{0,1}e_{0,\infty}-(1+(e_{0,1}/e_{1,\infty})+(e_{0,\infty}/e_{1,\infty})).$
@@
The result follows from (1). ∎
A cover $\phi$ can be described via the numerical data of the ramification type. However, there are also group-theoretic constraints on the ramification. To state these, we identify $G$ with $\mathrm{Gal}(\phi)$ and define the inertia type.
@@
Following Bouw [Bou01, Definitions 3.2 & 4.1], the inertia type of $\phi$ is the triple $a=(a_{0},a_{1},a_{\infty})$, where $a_{b}$ denotes the canonical generator for $I_{b}$, found from the action of inertia on a local parameter at a ramification point above $b$.
There are three constraints on the inertia type: $|a_{b}|=c_{b}$ for $b\in B$; $a_{0},a_{1},a_{\infty}$ generate $G$; and $a_{0}a_{1}a_{\infty}=\mathrm{Id}_{G}$. This implies that $s$ divides $c_{1}$ and $s$ divides $c_{\infty}$. Given a tuple $z=[c_{0},c_{1},c_{\infty},s]$ satisfying the numerical constraints for a ramification type, it is apriori possible that there is no tuple $a$ satisfying the group-theoretic constraints for the inertia type. See Remark 5.1 for some examples.
Given an inertia type, we can change the labeling of branch points whose inertia groups have the same order, permuting those entries of the inertia type. Also, we can change the identification of $G$ with $\mathrm{Gal}(\phi)$; this modifies each entry of the inertia type by some $\sigma\in\mathrm{Aut}(G)$. Two inertia types $a$ and $a^{\prime}$ are equivalent, denoted $a\approx a^{\prime}$, if they are in the same orbit under these actions.
@@ 2.2. Cyclic covers branched at three points
Suppose $\phi:X\to{\mathbf{P}}^{1}$ is cyclic of order $m$, and identify $G=\mathbf{Z}/m\mathbf{Z}$. The inertia type $(a_{0},a_{1},a_{\infty})$ is defined up to multiplication by $(\mathbf{Z}/m\mathbf{Z})^{*}$. Note that $c_{b}=|a_{b}|=m/\gcd(m,a_{b})$. An affine equation for the cover is
(3) $X:y^{m}=x^{a_{0}}(1-x)^{a_{1}},$
@@
2.3. Newton polygons
If $X$ is a smooth curve over $k$, we describe the Newton polygon of its Jacobian ${\mathbf{Jac}}(X)$ in terms of the decomposition of the $p$-divisible group ${\mathbf{Jac}}(X)[p^{\infty}]$.
Definition 2.6.
Suppose $A$ is a principally polarized abelian variety of dimension $g$ over $k$. By the Dieudonné–Manin classification [Man63], the $p$-divisible group $A[p^{\infty}]$ is isogenous to $\oplus_{\lambda}G_{\lambda}^{m_{\lambda}}$ for some multi-set of non-negative integers $m_{\lambda}$. This multi-set defines a symmetric Newton polygon of height $2g$ with integer break-points. As a short-hand, we write $\nu(A)=\oplus_{\lambda}G_{\lambda}^{m_{\lambda}}$ and call both $\nu(A)$ and its visual representation the Newton polygon of $A$.
The Newton polygon of a curve $X$ over $k$ is the Newton polygon of ${\mathbf{Jac}}(X)$. A curve $X$ of genus $g$ is ordinary if $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{g}$ and is supersingular if $\nu({\mathbf{Jac}}(X))=G_{1,1}^{g}$.
2.4. Newton polygons for cyclic covers branched at three points
Suppose $\phi:X\to\mathbf{P}^{1}$ is a cyclic degree $m$ cover which is branched at three points. Assuming that $p$ does not divide $m$, we describe how to compute the Newton polygon of ${\mathbf{Jac}}(X)$ from the signature type of $\phi$, using a formula due to Shimura and Taniyama, see [LMPT19b, §2,3] for details. This implicitly relies on $X$ being the reduction of a curve over $\mathbf{Q}$ whose Jacobian has complex multiplication.
Notation 2.7.
Assume that $p$ does not divide $m$. Let $R$ denote the set of orbits of the multiplication by $p$ map [×$[\times p]$ on the set $\{j\in\mathbf{Z}/m\mathbf{Z}:j\neq 0\}$. For each orbit $r\in R$, the additive order $d_{j}$ of $j$ modulo $m$ is the same for all values $j$ in $r$ and we denote it by $d_{r}$. For $\epsilon\in\{0,1\}$, let
[…]
In the context of Notation 2.7, the set $R$ is in one-to-one correspondence with the set of primes of the group ring $\mathbf{Q}[\mu_{m}]$ above $p$. This is because $\mathbf{Q}[\mu_{m}]$ is unramified at $p$; the elements of $\mathbf{Z}/m\mathbf{Z}$ can be identified with homomorphisms from $\mathbf{Q}[\mu_{m}]$ to $\mathbf{C}$; and there is a natural action of Frobenius on the set of homomorphisms, see [LMPT19b, §2.1] for details. The $p$-divisible group ${\mathbf{Jac}}(X)[p^{\infty}]$ canonically decomposes into pieces ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ indexed by the orbits $r\in R$.
Theorem 2.8 (Shimura–Taniyama formula).
[Tat71, Section 5] The only slope of the Newton polygon of ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ is $\lambda_{r}:=\alpha_{r}/\#r$.
In particular, if every orbit is self-dual, meaning invariant under negation modulo $m$, then $X$ is supersingular.
Example 2.9.
If $a=(1,1,6)$, then $\phi$ has signature $f=(1,1,1,0,0,0,0)$. Thus $S_{1}=\{1,2,3\}$. Then $\lambda_{r_{1}}=1$, $\lambda_{r_{2}}=1/2$, and $\lambda_{r_{5}}=0$. Thus $\nu(\mathbf{Jac}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$.
Remark 2.10.
There is another method to determine whether $X$ is supersingular by Yui [Yui80] and Aoki [Aok08] using $p$-adic valuations of Jacobi sums. By [Aok08, Theorem 1.1], $X$ is supersingular if and only if either (i) $p^{i}\equiv-1\bmod m$ for some $i$; or (ii) there is an equivalence (as defined in Section 2.1) of inertia types $a\approx(1,-p^{i},p^{i}-1)$ for some $i$ such that $d=\mathrm{gcd}(p^{i}-1,m)>1$ and $p^{j}\equiv-1\bmod m/d$ for some $j$.
@@ 2.5. Non-cyclic Galois groups and inertia types
Suppose $\phi$ is abelian but not cyclic, of order $m$. The following lemma determines the Galois group of $\phi$ in all the cases with $s>1$ for $5\leq g\leq 10$.
@@
By hypothesis, there is a prime $\ell$ dividing all three of $s=c_{0}s/c_{0}$, $c_{0}s/c_{1}$, and $c_{0}s/c_{\infty}$. If $G$ was cyclic, $\ell$ would divide each of $m=c_{0}s,a_{0},a_{1},a_{\infty}$, contradicting the fact that $a_{0},a_{1},a_{\infty}$ generate $G$. Since $G$ is generated by two elements $a_{0},a_{1}$, this implies that $G\simeq\mathbf{Z}/r_{1}\mathbf{Z}\times\mathbf{Z}/r_{2}\mathbf{Z}$ where $r_{1}r_{2}=c_{0}s$ and $r_{2}\mid r_{1}$. Note that $r_{1}$ is a multiple of $c_{0}$ since $a_{0}$ has order $c_{0}$. The hypothesis on $z$ implies that the orders of $a_{1}$ and $a_{\infty}$ divide $c_{0}$. So $a_{0}$ and $a_{1}$ generate $G$ only if $r_{1}=c_{0}$. Thus $G\simeq\mathbf{Z}/c_{0}\mathbf{Z}\times\mathbf{Z}/s\mathbf{Z}$. ∎
For $s>1$, we denote elements of $G=\mathbf{Z}/c_{0}\mathbf{Z}\times\mathbf{Z}/s\mathbf{Z}$ as an ordered pair. Each $a_{b}$ in the inertia type $a=(a_{0},a_{1},a_{\infty})$ is written as $a_{b}=(a_{b,1},a_{b,2})$. To choose a representative for the equivalence class of $a$, we take the natural inclusion of $G$ in $V=(\mathbf{Z}/c_{0}\mathbf{Z})^{2}$. Since $\mathrm{Aut}(V)$ acts transitively on pairs of linearly independent elements of order $c_{0}$, we may choose $a_{0}=(1,0)$ without loss of generality. When $c_{\infty}=s$, we also choose $a_{\infty}=(0,1)$.